# Knowledge in multi-robot systems: an interplay of dynamics, computation and communication

### v1

@@ 3.4 System Runs and Epistemic Frames
The remaining of this paper will focus on the computational aspect of the distributed multi-robot systems described in Section 3.3, i.e., on their epistemic states and on the transitions between epistemic states. The main tool that we are going to use to describe the epistemic evolution of robotic systems are epistemic frames, that we build from the system runs of Definition 3.5:
@@
Given a set of robots $\Pi$ and the set of all possible runs $\Sigma$, an epistemic frame =$\mathcal{F}=\langle\{\mathcal{C}\}_{\mathcal{C}\in\Sigma},\{\sim_{\mathbf{r}}\}_{\mathbf{r}\in\Pi}\rangle$ is composed of a set of global robot configurations $\{\mathcal{C}\}_{\mathcal{C}\in\Sigma}$ (Definition 3.4) and a collection of indistinguishability relations $\{\sim_{\mathbf{r}}\}_{\mathbf{r}\in\Pi}$, where:
  • Each global robot configuration, or point, is constituted by a pair $(\sigma,[t])$ of system run $\sigma$ (Definition 3.5) and the corresponding global timestamp $[t]$ (Definition 3.1), which is relative to the run $\sigma\in\Sigma$;
@@
The structure of the epistemic frame, and in particular its indistinguishability relations, depends heavily on the assumptions of the robot model (Table 1). To give some examples, if one assumes non-rigid movement then the robots will have greater uncertainty concerning their current location; visibility influences how much information can be gained during a look step; asynchronicity allows indistinguisbility to range among points with exact same global state but different global time, and so on.
4 Robot Tasks and Temporal Epistemic Logic
In this section we use the obtained epistemic frames to reason epistemically about robot tasks. We start by formalizing exploration tasks, and then we extend and refine this framework so as to accommodate the specifics of surveillance and approximate gathering tasks.
In robot tasks, we assume that robots are able to sense a portion of the environment in which they are embedded, typically represented as a topological space. We call the “observable” portion of the environment the exploration space, which we assume to be a compact topological space, denoted by $(X,\tau)$. Recall that a pair $(X,\tau)$ is a topological space if $X$ is a set, $\tau\subseteq 2^{X}$, $\varnothing\in\tau$, $X\in\tau$, and $\tau$ is closed under arbitrary unions, and finite intersections. $X$ is usually called the “space”, and $\tau$ is the collection of open sets. Recall that $(X,\tau)$ is compact iff for any $S\subseteq 2^{\tau}$ such that $\bigcup_{U\in S}U=X$, then there exists a finite subset $S^{\prime}\subseteq S$ such that $\bigcup_{U\in S^{\prime}}U=X$. Recall that a set $V\subseteq X$ is closed if $X\setminus V\in\tau$. For simplicity and illustrative purposes, we will assume that the exploration space is a compact normalized Euclidean space [^2] The Heine-Borel theorem states that in Euclidean spaces, compactness is equivalent to the space being bounded and closed in $\mathbb{R}^{n}$., i.e. $X\subseteq[0,1]^{n}$ with the standard Euclidean metric. Recall that in a metric space, a set $U\subseteq X$ is open iff for any $x\in U$, there exists an $\varepsilon>0$ such that $B_{\varepsilon}(x):=\{y\in X\;|\;d(x,y)<\varepsilon\}\subseteq U$.
Since we are interested in determining whether or not a region has been explored, we represent statements about exploration, such as “region $U$ has been explored”, using propositions: for every open set $U$ of $\tau$, we associate a unique proposition $\text{exp}(U)$, representing the fact that region $U$ is explored. $\text{exp}(X)$ expresses the fact that the whole space $X$ is explored, and $\text{exp}(\varnothing)$ represents that no portion of $X$ is explored. Because of the natural correlation between proposition and explored space, we refer to $\text{P}_{X}$ as the set of exploration statements.
Exploration statements satisfy some important properties: $V\supseteq U$ implies that $\text{exp}(V)\rightarrow\text{exp}(U)$. Thus, $(\text{P}_{X},\rightarrow)$ is a poset, induced by the containment of open sets. Furthermore, since $(X,\tau)$ is a topological space, and thus closed under union, then $(\text{P}_{X},\rightarrow)$ is a join-lattice. This property allows us to express exploration in a compact way, namely, by considering the largest explored region instead of considering separately all the possible smaller regions that have been explored. For instance, a robot a need only communicate to another robot b, the largest possible region that it has explored.
4.1 A Temporal Epistemic Logic for Exploration
@@
One main advantage of epistemic frames is that we can interpret indistinguishability relations as capturing the robots’ knowledge. Intuitively, a robot $\mathbf{r}$ knows a formula $\psi$ at $(\sigma,[t])$ iff such a formula holds at all points that are indistinguishable from $(\sigma,[t])$ to $\mathbf{r}$. In particular, because we defined indistinguishability between points on equivalence of robots’ local states, an S5 modal logic [^3] In an S5 modal logic, the knowledge operator satisfies factivity ($K_{a}\psi\rightarrow\psi$), positive introspection ($K_{a}\psi\rightarrow K_{a}K_{a}\psi$) and negative introspection ($\neg K_{a}\psi\rightarrow K_{a}\neg K_{a}\psi$). is the natural candidate. This notion of knowledge can be extended to groups, such as distributed knowledge via the indistinguishability relation $\sim_{\mathcal{D}_{A}}:=\bigcap_{\mathbf{r}\in A}\sim_{\mathbf{r}}$ consisting of the intersection of indistinguishabilities. It is easy to verify that $\sim_{\mathcal{D}_{A}}$ is also an equivalence relation. Intuitively, distributed knowledge represents the “collective knowledge” that is spread across all robots in $A$. Another important notion of group knowledge is mutual knowledge, which corresponds to everybody in a group knows a given formula $\psi$.
An epistemic frame can be enriched to an epistemic model by adding an evaluation function that assigns a truth value to atoms at each point of the model [^4] Since we focus mainly on exploration tasks, we restrict our attention to exploration statements, $\text{P}_{X}$. Nevertheless, note that the set of propositions may be chosen otherwise, and which one suits the best depends on the task at hand..
@@
Formally, an epistemic model $\mathcal{M}$ is a tuple of an epistemic frame $\mathcal{F}$ and an evaluation function $\pi:\text{P}_{X}\rightarrow 2^{\{\mathcal{C}\}_{\mathcal{C}\in\Sigma}}$, where $\mathcal{C}$ is a global robot configuration, and $\text{P}_{X}$ is the collection of exploration statements. In our particular case, the function $\pi$ simply maps a proposition $\text{exp}(U)$ corresponding to a region $U$, to all the global configurations where $U$ has been jointly explored (with or without communication) by the robots. In line with the runs and systems framework, we call our epistemic model interpreted system, denoted by $\mathcal{I}$.
Remark 4.1.
Note that, in the interpreted system usually found in the literature, a ”full-information“ protocol is used as a simple protocol that is comparable to the best obtainable one. Such protocol is not possible to be found for the case of distributed multi-robot systems, as robots necessarily need to make mutually exclusive choices, such as which direction to go. Therefore we fix a protocol, in the function $\varphi$ of Definition 3.6, and consider that indistinguishability in our runs ranges over the possible adversarial behavior of the environment.
The following grammar defines temporal epistemic logic formulas for exploration:
Definition 4.2 (Language $\mathcal{L}_{exp}$).
The exploration language is defined by the following grammar:
$\varphi::=\text{exp}(U)\>|\>\neg\varphi\>|\>(\varphi\wedge\varphi)\>|\>K_{\mathbf{r}}\varphi\>|\>D_{A}\varphi\>|\>\Diamond\varphi$
where $\text{exp}(U)\in\text{P}_{X}$, $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$.
@@
In this grammar, $\text{exp}(U)$ is an exploration statement, $K_{\mathbf{r}}\varphi$ represents “$\mathbf{r}$ knows $\varphi$”, $D_{A}\varphi$ stands for “ group A has distributed knowledge of $\varphi$”, and $\Diamond\varphi$ reads “eventually $\varphi$”.
Definition 4.3 (Semantics of $\mathcal{L}_{exp}$).
@@
Let $\mathcal{I}$ be a set of runs, with its respective set of configuration-time pairs $\mathcal{W}$, an evaluation function $\pi:\text{P}_{X}\rightarrow 2^{\mathcal{W}}$, a run $\sigma\in\mathcal{I}$, and a robot $\mathbf{r}\in\Pi$, :
  • $\mathcal{I},(\sigma,[t])\vDash\text{exp}(U)\text{ iff }\text{exp}(U)\in\pi(\text{exp}(U))$
  • $\mathcal{I},(\sigma,[t])\vDash\neg\varphi\text{ iff }\mathcal{I},(\sigma,[t])\not\vDash\varphi$
  • $\mathcal{I},(\sigma,[t])\vDash(\varphi\wedge\psi)\text{ iff }\mathcal{I},(\sigma,[t])\vDash\varphi\text{ and }\mathcal{I},(\sigma,[t])\vDash\psi$
@@
  • $\mathcal{I},(\sigma,[t])\vDash K_{\mathbf{r}}\varphi\text{ iff }\mathcal{I},(\sigma^{\prime},[t]^{\prime})\vDash\varphi\text{ for all }(\sigma^{\prime},[t]^{\prime})\in\mathcal{W}\text{ s.t. }(\sigma,[t])\underset{\mathbf{r}}{\sim}(\sigma^{\prime},[t]^{\prime})$
  • $\mathcal{I},(\sigma,[t])\vDash D_{A}\varphi\text{ iff }\mathcal{I},(\sigma^{\prime},[t]^{\prime})\vDash\varphi\text{ for all }(\sigma^{\prime},[t]^{\prime})\in\mathcal{W}\text{ s.t. }(\sigma,[t])\underset{{\mathcal{D}_{A}}}{\sim}(\sigma^{\prime},[t]^{\prime})$
@@
  • $\mathcal{I},(\sigma,[t])\vDash\Diamond\varphi$ iff $\mathcal{I},(\sigma,[t]^{\prime})\vDash\varphi$ for some $[t]^{\prime}\geqslant[t]$.
@@
Parenthesis are omitted in favor of readability whenever precedence is not ambiguous.
A formula $\varphi$ is valid in $\mathcal{I}$ if, for all $(\sigma,[t])\in\mathcal{W}$, $\mathcal{I},(\sigma,[t])\vDash\varphi$. If $\varphi$ is valid in $\mathcal{I}$, we use the notation $\mathcal{I}\vDash\varphi$.
Remark 4.4.
If an robot knows that a certain region $\text{exp}(U)$ has been explored, formally $K_{\mathbf{r}}\text{exp}(U)$, it implies that all subsets of that region are also known by that robot, i.e., $K_{\mathbf{r}}\text{exp}(V)$ with $V\subseteq U$ and $V$ open. This follows directly from the compactness of the exploration space.
Having established the temporal epistemic framework, it is evident that the space will be explored at a run $\sigma\in\mathcal{I}$ iff $\mathcal{I},\sigma\vDash\Diamond\text{exp}(X)$.
Although this condition seems simple enough to achieve, a more interesting part of the analysis lies within the different termination conditions for the exploration problem. In what follows, we introduce the epistemic characterization of the simple exploration robot task and different termination conditions for it.
@@
We say that a system is consistent with a simple exploration system if it satisfies:
  • Exploration agency:
    $\mathcal{I}\vDash\text{exp}(U)\rightarrow D_{\Pi}\text{exp}(U)$ (16)
  • Exploration independence:
    $\mathcal{I}\vDash D_{A}\text{exp}(U)\rightarrow\bigwedge_{\mathbf{r}\in A}K_{\mathbf{r}}exp(V_{\mathbf{r}})\textrm{ such that }\bigcup_{\mathbf{r}\in A}V_{\mathbf{r}}=U$ (17)
  • Stable exploration statements:
    For any region $U$
    $\mathcal{I}\vDash\bigwedge_{\text{exp}(U)\in\mathcal{L}_{exp}^{-}}\text{exp}(U)\rightarrow\Box\text{exp}(U)$ (18)
  • Stable exploration knowledge (perfect recall):
    For any $\mathbf{r}\in\Pi$, and any region $U$
    $\mathcal{I}\vDash K_{\mathbf{r}}\text{exp}(U)\rightarrow\Box K_{\mathbf{r}}\text{exp}(U)$ (19)
Exploration agency claims that an atom is true, meaning that a region is explored, because it has been explored by at least one robot in the group.
Exploration independence means that whatever exploration atom the robots know distributively, is given by the union of the exploration atoms known by those robots. In other words, exploration (distributed) knowledge can be acquired either by direct observation, or robot communication and not by other inferences about the system.
Stable exploration statements states that once a region has been explored, it will be remain forever explored. Stable exploration knowledge says that once an exploration statement is known, it will be known forever by that robot.
The following definitions formally state different notions of termination that may be considered for the exploration task, and possibly for robot tasks in general.
@@
Given an interpreted system $\mathcal{I}$ that is consistent with simple exploration, a set of robots $\Pi$, with an exploration space $X$, and its respective exploration statements $\text{P}_{X}$, we define the following termination conditions:
  • Parallel Termination: We say that $\mathcal{I}$ is consistent with parallel termination iff $\mathcal{I}\vDash\Diamond D_{\Pi}\text{exp}(X)$.
  • Selfish Termination: We say that $\mathcal{I}$ is consistent with selfish termination iff $\mathcal{I}\vDash\Diamond E_{\Pi}\text{exp}(X)$.
  • Cooperative Termination: We say that $\mathcal{I}$ is consistent with cooperative termination iff $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$.
Intuitively, the parallel termination condition corresponds to a model where a global control or observer is able to terminate the task whenever it is completed, the selfish termination condition corresponds to a setting where it is sufficient for each robot to be aware that the exploration is complete in order to terminate, and cooperative termination corresponds to a setting where a robot may only terminate whenever it is certain that eventually all other processes will be aware that the space exploration is complete.
In the following, we will characterize parallel termination in terms of a very weak condition, which we call exploration liveness. Intuitively, exploration liveness means that any point $x\in X$ of the exploration space $X$ will eventually be explored. Note that, since we assume that robot exploration space consists only of open sets of $X$, the exploration liveness condition reduces to the existence of a (possibly infinite) open cover of $X$ that consists of explored regions, i.e., there must exist a collection $\mathcal{U}$ of open sets, such that $\bigcup_{U\in\mathcal{U}}U=X$, and eventually each $U$ is explored.
@@
We say that a run $\sigma$ satisfies exploration liveness iff there exists a collection of open sets, $\mathcal{U}$, such that $\bigcup_{U\in\mathcal{U}}U=X$, and for any $U\in\mathcal{U}$, $\mathcal{I},\sigma\vDash\Diamond\text{exp}(U)$. Respectively, we say that a model satisfies exploration liveness iff all of its runs satisfy exploration liveness.
@@
Proof 4.9.
Note by contraposition that if there is a run $\sigma$, such that no open cover $\mathcal{U}$ of eventually explored regions exists. Naturally, we may consider $\mathcal{U}^{\prime}=\{U\;|\;\mathcal{I},\sigma\vDash\Diamond\text{exp}(U)\}$. Since $\mathcal{U}^{\prime}$ is not an open cover of $X$, then $\bigcup_{U\in\mathcal{U}^{\prime}}U\neq X$. Therefore, there is a point $x\in X,x\notin\bigcup_{U\in\mathcal{U}^{\prime}}U$. Thus, by definition of $\mathcal{U}^{\prime}$, $x$ is never explored, and consequently $X$ is not explored in $\sigma$.
@@
Let $\mathcal{I}$ be a model with a compact exploration space $X$. If $\mathcal{I}$ satisfies exploration liveness, then $\mathcal{I}\vDash\Diamond\text{exp}(X)$.
Proof 4.11.
Consider an arbitrary run $\sigma$ of $\mathcal{I}$, from Definition 4.7, there is an open cover $\mathcal{U}$ such that for any $U\in\mathcal{U}$, $\mathcal{I},\sigma\vDash\Diamond\text{exp}(U)$. Since $X$ is compact and $\mathcal{U}$ is an open cover of $X$ then there exists a finite subcover $\mathcal{U}^{\prime}=\{V_{1},\ldots,V_{k}\}\subseteq\mathcal{U}$, such that $\bigcup_{i=1}^{k}V_{i}=X$. Also, since $\mathcal{I},\sigma\vDash\Diamond\text{exp}(V_{i})$, then for any $i\in[1,k]$, then for each $V_{i}$ there is a time $[t_{i}]$ such that $\mathcal{I},\sigma,[t_{i}]\vDash\text{exp}(V_{i})$. Simply consider $[t_{\textrm{max}}]=\mbox{}_{\textrm{max}}\{[t_{i}]\}$. Since exploration statements are stable facts, then $\mathcal{I},\sigma,[t_{\textrm{max}}]\vDash\bigwedge_{i=1}^{k}\text{exp}(V_{i})$. Therefore $\mathcal{I},\sigma,[t_{\textrm{max}}]\vDash\text{exp}(X)$. Thus $\mathcal{I},\sigma\vDash\Diamond\text{exp}(X)$. Since $\sigma$ is an arbitrary run of $\mathcal{I}$, then it follows that $\mathcal{I}\vDash\Diamond\text{exp}(X)$.
Note that Theorem 4.8, Theorem 4.10, and exploration agency from Definition 4.5 imply that whenever $X$ is compact, then exploration liveness is necessary and sufficient for parallel termination. However this is not necessarily true in non-compact spaces. For instance, consider a single robot exploring the open interval $(0,1)\subset\mathbb{R}$. Consider a scenario where at each step $[t]$ the robot has explored the region $(0,1-(1/2)^{[t]})$. It is easy to verify that in this case, the exploration liveness condition is satisfied, nevertheless $\Diamond\text{exp}(X)$ never holds, since at any given time, there is still a portion of the space that remains to be explored, namely $[1-(1/2)^{[t]},1)$.
4.2 The Exploration Task in the LCM Model
We first consider a $\mathcal{LUMI}$ robot system, as defined in Definition 2.7. Recall that robots communicate by displaying information via a set of lights that is assumed to remain active at all times, and that in the full-visibility case is observable at any time by an active robot in the look phase of its LCM cycle.
@@ 4.2 The Exploration Task in the LCM Model
We say that the communication of a robot system is exploration flooding iff at each phase of the LCM cycle, each robot tries to communicate the largest region it knows to be explored to the rest of the robots.

### v2

@@ 4 Robot Tasks and Temporal Epistemic Logic
The remaining of this paper focuses on the computations of the distributed multi-robot systems described in Section˜3, i.e., on their epistemic states and on their transitions. The epistemic evolution of robotic systems is described using epistemic frames that are build from the system runs of Definition˜3.5:
@@
Given a set of robots $\Pi$ and the set of all possible runs $\Sigma$, an epistemic frame $\mathcal{F}=\langle\{\mathcal{C}\}_{\mathcal{C}\in\Sigma},\{\sim_{\mathbf{r}}\}_{\mathbf{r}\in\Pi}\rangle$ is composed of a set of global robot configurations $\{\mathcal{C}\}_{\mathcal{C}\in\Sigma}$ (Definition˜3.4) and a collection of indistinguishability relations $\{\sim_{\mathbf{r}}\}_{\mathbf{r}\in\Pi}$, where:
  • Each global robot configuration, or point, is constituted by a pair $(\sigma,[t])$ of system run $\sigma$ (Definition˜3.5) and the corresponding global timestamp $[t]$ (Definition˜3.1), which is relative to the run $\sigma\in\Sigma$;
@@
The structure of the epistemic frame, and in particular its indistinguishability relations, depends on the assumptions of the robot model. For instance, if one considers that robots may not complete their move steps (non-rigid model), uncertainty regarding their current locations will be added; visibility influences how much information can be gained (and discerned) during a look step; asynchronicity allows indistinguisbility to range among points with the exact same global state but different global time, and so on.
The epistemic frames, generated from the system runs of Definition˜3.5, are used to reason about the knowledge of the multi-robot system throughout the robot tasks. Fittingly, we consider only the epistemic states of the system, as the ontic (physical) states are abstracted away in the transitions between computation steps.
In robot tasks, we assume that robots are able to sense a portion of the environment in which they are embedded, represented as a topological space. We call the “observable” portion of the environment the observable space, which we assume to be a compact topological space, denoted by $(X,\tau)$. In particular, it is a compact normalized Euclidean space, i.e. $X\subseteq[0,1]^{n}$ with the standard Euclidean metric. To every open set $U$ of $\tau$ we associate a unique proposition $\text{sp}(U)$. $\text{sp}(X)$ is associated to the whole observation space $X$ and $\text{sp}(\varnothing)$ to the empty space $\varnothing$. We refer to $\text{P}_{X}$ as the set of space statements. Space statements satisfy some important properties: $V\supseteq U$ implies that $\text{sp}(V)\rightarrow\text{sp}(U)$. Thus, $(\text{P}_{X},\rightarrow)$ is a poset, induced by the containment of open sets. Furthermore, since $(X,\tau)$ is closed under union $(\text{P}_{X},\rightarrow)$ is a join-lattice, allowing us to express, for example, information in a compact way: robot a need only communicate to another robot b, the largest region that contains all others, i.e., a single atom.
@@
We say that a robot $\mathbf{r}$ knows a formula $\psi$ at $(\sigma,[t])$ iff such a formula holds at all points that are indistinguishable from $(\sigma,[t])$ to $\mathbf{r}$. In particular, because indistinguishability between points is defined on equivalence of robots’ local states, the S5 epistemic logic [^1] In an S5 modal logic, the knowledge operator satisfies factivity ($K_{a}\psi\rightarrow\psi$), positive introspection ($K_{a}\psi\rightarrow K_{a}K_{a}\psi$) and negative introspection ($\neg K_{a}\psi\rightarrow K_{a}\neg K_{a}\psi$). is the natural candidate. This notion of knowledge can be extended to groups, such as distributed knowledge $(\mathcal{D}_{A}\varphi)$ via the equivalence relation $\sim_{\mathcal{D}_{A}}:=\bigcap_{\mathbf{r}\in A}\sim_{\mathbf{r}}$ consisting of the intersection of indistinguishabilities and representing the “collective knowledge” spread across all robots in $A$.
@@
An epistemic model $\mathcal{M}$ is a tuple of an epistemic frame $\mathcal{F}$ and an evaluation function $\pi:\text{P}_{X}\rightarrow 2^{\{\mathcal{C}\}_{\mathcal{C}\in\Sigma}}$ that assigns a truth value to atoms at each point of the model, where $\mathcal{C}$ is a global robot configuration, and $\text{P}_{X}$ is the collection of space statements The function $\pi$ simply maps a proposition $\text{sp}(U)$ corresponding to a region $U$, to all the global configurations where $U$ holds. In line with the runs and systems framework, we call our epistemic model interpreted system, denoted by $\mathcal{I}$.
Note that, even though our logic is factive, known facts about the real world are not necessarily true in the real world. As knowledge of robots depends on the robots observation, factivity can only hold with respect to that observation. Consider as an example the formula $(\sigma,[t])\models K_{\mathbf{r}}position(\mathbf{r},x,y)$ denoting that robot $\mathbf{r}$’s position at the time $t$ is at the point $(x,y)$ and lets assume that this was actually true in the actual world at the time robot observed its surrounding. This does not imply that it is still true at any time $t^{\prime}\neq t$ where $t^{\prime}]=[t]$. In fact our knowledge operator is only S5 with respect to the epistemic observations, but not with respect to the real world. Common knowledge experiences similar frictions, where attainability depends on the model and timing assumptions [16, 22], yet our modeling unveils the crucial dependency on observations. This is in line with the fact that we consider only the epistemic states of the system, abstracting away the ontic (physical) states in the transitions between computation steps.
4.1 A Temporal Epistemic Logic of Space
The following grammar defines temporal epistemic logic formulas:
Definition 4.2 (Language $\mathcal{L}_{space}$).
The language is defined by the following grammar:
$\varphi::=\text{sp}(U)\>|\>\neg\varphi\>|\>(\varphi\wedge\varphi)\>|\>K_{\mathbf{r}}\varphi\>|\>D_{A}\varphi\>|\>\Diamond\varphi$
where $U\in\text{P}_{X}$, $\mathbf{r}\in\Pi$, and $A\subseteq\Pi$.
@@ 4.1 A Temporal Epistemic Logic of Space
In this grammar, $\text{sp}(U)$ is a space statement, $K_{\mathbf{r}}\varphi$ represents $\mathbf{r}$ knows $\varphi$”, $D_{A}\varphi$ stands for “ group A has distributed knowledge of $\varphi$”, and $\Diamond\varphi$ reads “eventually $\varphi$”.
Definition 4.3 (Semantics of $\mathcal{L}_{space}$).
@@
Let $\mathcal{I}$ be a set of runs (Definition˜4.1), with its respective set of configuration-time pairs $\mathcal{W}$, an evaluation function $\pi:\text{P}_{X}\rightarrow 2^{\mathcal{W}}$, a run $\sigma\in\mathcal{I}$, and a robot $\mathbf{r}\in\Pi$, :
  • $\mathcal{I},(\sigma,[t])\models\text{sp}(U)\text{ iff }\text{sp}(U)\in\pi(\sigma,[t])(\text{sp}(U))$
[…]
  • $\mathcal{I},(\sigma,[t])\models(\varphi\wedge\psi)\text{ iff }\mathcal{I},(\sigma,[t])\models\varphi\text{ and }\mathcal{I},(\sigma,[t])\models\psi$
@@
  • $\mathcal{I},(\sigma,[t])\models K_{\mathbf{r}}\varphi\text{ iff }\mathcal{I},(\sigma^{\prime},[t]^{\prime})\models\varphi\text{ for all }(\sigma^{\prime},[t]^{\prime})\in\mathcal{W}\text{ s.t. }(\sigma,[t])\underset{\mathbf{r}}{\sim}(\sigma^{\prime},[t]^{\prime})$
  • $\mathcal{I},(\sigma,[t])\models D_{A}\varphi\text{ iff }\mathcal{I},(\sigma^{\prime},[t]^{\prime})\models\varphi\text{ for all }(\sigma^{\prime},[t]^{\prime})\in\mathcal{W}\text{ s.t. }(\sigma,[t])\underset{{\mathcal{D}_{A}}}{\sim}(\sigma^{\prime},[t]^{\prime})$
@@
  • $\mathcal{I},(\sigma,[t])\models\Diamond\varphi$ iff $\mathcal{I},(\sigma,[t]^{\prime})\models\varphi$ for some $t]^{\prime}\geqslant[t]$.
@@
A formula $\varphi$ is valid in $\mathcal{I}$ if, for all $(\sigma,[t])\in\mathcal{W}$, $\mathcal{I},(\sigma,[t])\models\varphi$. If $\varphi$ is valid in $\mathcal{I}$, we use the notation $\mathcal{I}\models\varphi$.
4.2 Application to Exploration
In this section we analyze the robot exploration task for a LCM model using the framework introduced in the previous section. Spatial atoms now denote that a certain region is explored, i.e., some robot has observed it, without changing the semantics.
@@ 4.2 Application to Exploration
We say that a system is consistent with a simple exploration system if, for any $\mathbf{r}\in\Pi$, and any region $U$, it satisfies:
  • Exploration agency: $\mathcal{I}\models\text{sp}(U)\rightarrow D_{\Pi}\text{sp}(U)$, i.e., if a region $U$ has been explored, it has been explored by at least one robot
  • Exploration independence: $\mathcal{I}\models D_{A}\text{sp}(U)\rightarrow\bigwedge_{\mathbf{r}\in A}K_{\mathbf{r}}\text{sp}({V_{\mathbf{r}}})\textrm{ s. t. }\bigcup_{\mathbf{r}\in A}V_{\mathbf{r}}=U$, i.e., any space atom known distributively by robots in $A$, is given by the union of the space atoms known by robots in AA
  • Stable space statements: $\mathcal{I}\models\bigwedge_{\text{sp}(U)\in\mathcal{L}_{exp}^{-}}\text{sp}(U)\rightarrow\Box\text{sp}(U)$, i.e., once a region has been explored, it is forever explored.
  • Stable exploration knowledge (perfect recall): $\mathcal{I}\models K_{\mathbf{r}}\text{sp}(U)\rightarrow\Box K_{\mathbf{r}}\text{sp}(U)$, i.e., once a space atom is known by a robot, it will be known forever by that robot.
In the exploration task, if a robot knows that a certain region of space $\text{sp}(U)$ has been explored, denoted by $K_{\mathbf{r}}\text{sp}(U)$, it implies that all subsets of that region are also known by that robot, i.e., $K_{\mathbf{r}}\text{sp}(V)$ with $V\subseteq U$ and $V$ open. This follows directly from the compactness of the exploration space.
@@
Given an interpreted system $\mathcal{I}$ consistent with simple robot exploration (Definition˜4.4), with a set of robots $\Pi$, exploration space $X$, and its respective space statements $\text{P}_{X}$, we say that $\mathcal{I}$ is consistent with:
  • Parallel termination: iff $\mathcal{I}\models\Diamond D_{\Pi}\text{sp}(X)$, i.e., no single robot needs to know that the whole space is explored;
  • Selfish termination: iff $\mathcal{I}\models\Diamond E_{\Pi}\text{sp}(X)$, i.e., each robot needs to know that the whole space is explored;
  • Cooperative termination: iff $\mathcal{I}\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{sp}(X)$, i.e., each robot needs to know that the whole space is explored and that other robots know that the whole space is explored.
Parallel termination can be characterized in terms of a very weak condition, called exploration liveness:
@@
A run $\sigma$ satisfies exploration liveness iff there exists a collection of open sets, $\mathcal{U}$, such that $\bigcup_{U\in\mathcal{U}}U=X$, and for any $U\in\mathcal{U}$, $\mathcal{I},\sigma\models\Diamond\text{sp}(U)$. A model satisfies exploration liveness iff all of its runs satisfy exploration liveness.
In other words, exploration liveness says that any point $x\in X$ of the exploration space $X$ will eventually be explored.
@@
Proof 4.8.
Note by contraposition that if there is a run $\sigma$, such that no open cover $\mathcal{U}$ of eventually explored regions exists. Naturally, we may consider $\mathcal{U}^{\prime}=\{U\;|\;\mathcal{I},\sigma\models\Diamond\text{sp}(U)\}$. Since $\mathcal{U}^{\prime}$ is not an open cover of $X$, then $\bigcup_{U\in\mathcal{U}^{\prime}}U\neq X$. Therefore, there is a point $x\in X,x\notin\bigcup_{U\in\mathcal{U}^{\prime}}U$. Thus, by definition of $\mathcal{U}^{\prime}$, $x$ is never explored, and consequently $X$ is not explored in $\sigma$.
@@
Let $\mathcal{I}$ be a model with a compact exploration space $X$. If $\mathcal{I}$ satisfies exploration liveness, then $\mathcal{I}\models\Diamond\text{sp}(X)$.
Proof 4.10.
Consider an arbitrary run $\sigma$ of $\mathcal{I}$, from Definition˜4.6, there is an open cover $\mathcal{U}$ such that for any $U\in\mathcal{U}$, $\mathcal{I},\sigma\models\Diamond\text{sp}(U)$. Since $X$ is compact and $\mathcal{U}$ is an open cover of $X$ then there exists a finite subcover $\mathcal{U}^{\prime}=\{V_{1},\ldots,V_{k}\}\subseteq\mathcal{U}$, such that $\bigcup_{i=1}^{k}V_{i}=X$. Also, since $\mathcal{I},\sigma\models\Diamond\text{sp}(V_{i})$, then for any $i\in[1,k]$, then for each $V_{i}$ there is a time $t_{i}]$ such that $\mathcal{I},\sigma,[t_{i}]\models\text{sp}(V_{i})$. Simply consider $t_{\textrm{max}}]=\mbox{}_{\textrm{max}}\{[t_{i}]\}$. Since space statements are stable facts, then $\mathcal{I},\sigma,[t_{\textrm{max}}]\models\bigwedge_{i=1}^{k}\text{sp}(V_{i})$. Therefore $\mathcal{I},\sigma,[t_{\textrm{max}}]\models\text{sp}(X)$. Thus $\mathcal{I},\sigma\models\Diamond\text{sp}(X)$. Since $\sigma$ is an arbitrary run of $\mathcal{I}$, then it follows that $\mathcal{I}\models\Diamond\text{sp}(X)$.
Theorem˜4.7, Theorem˜4.9, and exploration agency from Definition˜4.4 imply that whenever $X$ is compact exploration liveness is necessary and sufficient for parallel termination [^2] Note that this is not necessarily true in non-compact spaces..
@@
We say that the communication of a robot system is exploration flooding iff at each phase of the LCM cycle, each robot tries to communicate the largest region of space it knows to be explored to the rest of the robots.